# The minimal periodicity for integral bases of pure number fields

### v1

@@ 4.2. Counting and density of shape classes
$\frac{g}{d^{n}}\,\mathbf{1}_{\{g\mid r\}}=\prod_{p}\Bigl{(}p^{\min\{n\nu_{p}(d),\alpha_{p}\}-n\nu_{p}(d)}\ \mathbf{1}_{\{\min\{n\nu_{p}(d),\alpha_{p}\}\leq v_{p}(r)\}}\Bigr{)},$
with the convention $\alpha_{p}=0$ when $p\nmid q$. Since $\mu(p^{k})=0$ for $k\geq 2$, only $\nu_{p}(d)\in\{0,1\}$ contribute, and the local factor at each prime $p$ becomes
$\begin{cases}(1-\tfrac{1}{p^{n}}),&p\nmid q,\\[4.0pt] 1-\mathbf{1}_{\{v_{p}(r)\geq\alpha_{p}\}}\,p^{\alpha_{p}-n},&p^{\alpha}\parallel q.\end{cases}$
Multiplying over $p$ gives the stated main term. If $r$ is not nn–admissible then there exists $p^{\alpha}\parallel q$ with $\alpha\geq n$ and $v_{p}(r)\geq n$, forcing $p^{n}\mid a$ for every $a\equiv r\pmod{q}$ and thus $\mathcal{F}_{n}(X;q,r)\in\{0,1\}$. If $r$ is strictly nn–admissible, then for each $p^{\alpha}\parallel q$ one has $v_{p}(r)<\min\{\alpha,n\}$, whence the pp–factor equals $1$, and the uniform main term follows. ∎
Remark 4.7 (Non–monogeneity of positive density).
In degree $n$ with at least one prime $p\parallel n$, Proposition 4.1 yields a local obstruction governed by the single congruence $a^{p-1}\not\equiv 1\pmod{p^{2}}$ at $p$. Among units modulo $p^{2}$ this occurs with proportion $1-\tfrac{1}{p}$. Combining Theorem 4.6 with the obvious squarefreeness constraints gives explicit positive lower bounds for the density (in height) of non–monogenic $K_{a}$ in any fixed degree $n$.
Corollary 4.8 (Uniformity on “nn–full” moduli).
@@
factors through $a\bmod M(n)$. Fix any union $\mathcal{R}\subseteq(\mathbb{Z}/M(n)\mathbb{Z})$ of admissible classes (i.e. classes $r$ with $v_{p}(r)\leq 1$ for each $p\mid n$). Let
@@
Write the indicator of $n$th–power–freeness as $1_{\text{$n$–free}}(a)=\sum_{d^{n}\mid a}\mu(d)$ and count by inclusion–exclusion in the progression $a\equiv r\pmod{M(n)}$. If $r$ is admissible, then $(d,M(n))=1$ in all contributing terms; the main term is
@@
Summing over $r\in\mathcal{R}$ gives the formula; standard partial summation bounds yield $o(X)$. If $r$ is not admissible there are $O(1)$ terms. ∎
@@
Here $M(4)=8$ and the admissible residue classes (in the sense after Theorem 4.10) are those r(mod 8)r\ ($\mathrm${mod}\ 8) with $v_{2}(r)\leq 1$, i.e. $r\in\{1,3,5,7,2,6\}$ (six classes). The three quartic shapes are:
@@
By Theorem 4.10, for any union $\mathcal{R}$ of admissible classes one has
@@
(These four densities sum to $\tfrac{24}{36}C_{6}$, reflecting that we have restricted to the 2424 admissible classes.)
@@
$(\mathbb{Z}/p^{t}\mathbb{Z})^{\times}$ is cyclic of order $p^{t-1}(p-1)$. The congruence $x^{p-1}\equiv 1\pmod{p^{t}}$ has exactly p1p-1 solutions and solutions lift uniquely with $t$ by Hensel’s lemma. Among $p^{e}(p-1)$ units modulo $p^{e+1}$, those with $v_{p}(u^{p-1}-1)\geq k$ are $(p-1)p^{e-(k-1)}$, giving the stated probabilities; the formulae follow. The case $p=2$ is similar with the usual structure of $(\mathbb{Z}/2^{t}\mathbb{Z})^{\times}$. ∎
@@
Appendix A The Ore–Montes toolkit for $f(x)=x^{n}-a$ at a fixed prime $p$
In this section we give a concise, self–contained exposition of the Ore–Montes (higher Newton polygon) method as it applies to the binomials $f(x)=x^{n}-a\in\mathbb{Z}[x]$, that play main role in the proof the minimal local (hence global) modulus controlling the shape of integral bases.
Fix a prime $p$, write $v=v_{p}$ for the normalized pp–adic valuation on $\mathbb{Q}_{p}$, and consider $f(x)=x^{n}-a$ with $a\in\mathbb{Z}$ $n$th–power–free. Let $p^{e}\parallel n$ ($e\geq 1$).
(1) Order 1 polygons and residual polynomials. Choose a monic lift $\varphi_{1}\in\mathbb{Z}_{p}[x]$ of an irreducible factor of $f(x)\bmod p$:
$\varphi_{1}=\begin{cases}x-u,&\text{if }v(a)=0\text{ with }u^{n}\equiv a\pmod{p},\\ x,&\text{if }v(a)>0.\end{cases}$
Write the $\varphi_{1}$–adic expansion $f=\sum_{i\geq 0}a_{1,i}(x)\,\varphi_{1}(x)^{i}$ with $a_{1,i}\in\mathbb{Z}_{p}[x]$ and $\deg a_{1,i}<\deg\varphi_{1}$, and let $u_{1,i}:=v(a_{1,i})$. The (principal) Newton polygon $N^{+}_{\varphi_{1}}(f)$ is the lower convex hull of the points $(i,u_{1,i})$ with $a_{1,i}\neq 0$. For each side $S$ of slope $\lambda_{1}=-h_{1}/e_{1}$ (coprime $h_{1},e_{1}\in\mathbb{Z}_{>0}$) whose lattice points are $(s,u_{1,s}),(s+e_{1},u_{1,s}-h_{1}),\dots,(s+d_{1}e_{1},u_{1,s}-d_{1}h_{1})$, the residual polynomial in $\mathbb{F}_{\varphi_{1}}:=\mathbb{F}_{p}[x]/(\overline{\varphi_{1}})$ is
$R_{1}(y)\ =\ \sum_{j=0}^{d_{1}}c_{1,j}y^{j},\qquad c_{1,j}\ :=\ \Bigl{(}\,\frac{a_{1,s+je_{1}}(x)}{p^{\,u_{1,s}-jh_{1}}}\,\Bigr{)}\ \bmod(p,\varphi_{1}).$
If each $R_{1}$ is separable, $f$ is pp–regular and order 1 already determines the pp–adic factorization and a pp–integral basis via Ore’s index theorem; otherwise pass to higher orders.
(2) Higher orders, types, and augmented valuations. A type of order t1t-1 is a datum
$\mathbf{t}_{t-1}=(\varphi_{1},-\lambda_{1};\,\varphi_{2},-\lambda_{2};\,\dots;\,\varphi_{t-1},-\lambda_{t-1};\,\psi_{t-1})$
where each $\varphi_{i}\in\mathbb{Z}_{p}[x]$ is a key polynomial, $\lambda_{i}=h_{i}/e_{i}>0$ is a slope, and $\psi_{t-1}\in\mathbb{F}_{t-1}[y]$ is the last residual polynomial over a finite residue field $\mathbb{F}_{t-1}$. The associated augmented valuation $\omega_{i+1}=e_{i}[\omega_{i},\varphi_{i},\lambda_{i}]$ on $\mathbb{Q}_{p}(x)$ is given on $\varphi_{i}$–adic expansions $F=\sum_{j}a_{i,j}\varphi_{i}^{j}$ by
$\omega_{i+1}(F)\ =\ \min_{j}\Bigl{(}e_{i}\,\omega_{i}(a_{i,j})+j\bigl{(}e_{i}\omega_{i}(\varphi_{i})+h_{i}\bigr{)}\Bigr{)}.$
At order $t$ one forms the $\varphi_{t}$–adic expansion with respect to $\omega_{t}$, builds the polygon $N_{t}(f)$, reads the residual polynomial $R_{t}(y)\in\mathbb{F}_{t}[y]$, and refines the type. If the residual at the last principal side is separable, the process stops and Ore’s theorem gives $\nu_{p}((\mathcal{O}_{K}:\mathbb{Z}[\theta]))$ together with a pp–integral basis.
(3) Special features for $f=x^{n}-a$ with $p^{e}\parallel n$ and $(a,p)=1$.
In the pure binomial setting the chain of types never exceeds order $e$, and the principal sides that actually occur have height $1$. Concretely, when $v(a)=0$ and $\varphi_{1}=x-u$, the order 1 polygon is a broken line with consecutive drops of height $1$ (as encoded by $v_{p}(a^{p-1}-1)$); when $v(a)>0$ and $\varphi_{1}=x$, the unique principal side again has height $1$. In either branch, each lift to order $t$ adds exactly one unit of height.
We isolate the bookkeeping that drives our local modulus.
Lemma A.1 (Kummer valuations for the unit branch).
Let $p^{e}\,\|\,n$ and $(a,p)=1$. In the $\varphi_{1}=x-u$ development with $u^{n}\equiv a\pmod{p}$,
$v_{p}\!\big{(}\tbinom{n}{i}u^{\,n-i}\big{)}\ \geq\ e-v_{p}(i)\qquad(1\leq i\leq n),$
with equality for $i=p^{k}$ ($0\leq k\leq e$). In particular $N^{-}_{\varphi_{1}}(f)$ contains the points $(p^{k},e-k)$, and each principal side joining $(p^{k-1},e-(k-1))$ to $(p^{k},e-k)$ has height $1$.
Proof.
By Kummer’s theorem, $v_{p}\binom{n}{i}$ equals the number of carries when adding $i$ and nin-i in base $p$. Since $p^{e}\,\|\,n$, one obtains $v_{p}\binom{n}{i}\geq e-v_{p}(i)$ with equality for $i=p^{k}$. Multiplying by the unit $u^{n-i}$ does not change the valuation. ∎
Lemma A.2 (Order–$t$ residuals use $a\bmod p^{t+1}$).
Let $p^{e}\parallel n$ and choose $\varphi_{1}$ as above. For each $t$ with $1\leq t\leq e$, the residual polynomial $R_{t}(y)$ built on a principal side of $N_{t}(f)$ depends only on the class of $a\bmod p^{t+1}$.
Proof of Lemma A.2.
By Lemma A.1, every principal side of $N_{t}(f)$ that actually occurs up to order $e$ has height $1$. Along such a side the residual coefficients have the form
$c_{t,j}\;=\;\overline{\,a_{t,s+je}/p^{\,u_{t,s}-j}\,}\in\mathbb{F}_{t},$
so they are determined by the coefficients modulo one more power of $p$ than at order t1t-1. Induction on $t$ shows that $R_{t}(f)$ depends only on $a\bmod p^{t+1}$. ∎
As consequence, if $p^{e}\parallel n$, then the entire pp–type of $f(x)=x^{n}-a$ (slopes, residual polynomials, and hence the pp–integral basis output by Ore–Montes) is determined by $a\bmod p^{e+1}$.

### v2

@@ 4.2. Counting and density of shape classes
$S(q,r)=\prod_{p\nmid q}(1-p^{-n})\cdot\prod_{p^{\alpha}\parallel q}\Bigl(1-\mathbf{1}_{\{v_{p}(r)\geq\min(\alpha,n)\}}\,p^{\min(\alpha,n)-n}\Bigr).$
Since $\prod_{p}(1-p^{-n})=\zeta(n)^{-1}$, we may rewrite
$\prod_{p\nmid q}(1-p^{-n})=\frac{1}{\zeta(n)}\prod_{p\mid q}(1-p^{-n})^{-1}.$
Substituting back gives precisely the main term in (3).
If $r$ is not $n$-admissible, then there exists $p^{\alpha}\parallel q$ with $\alpha\geq n$ and $v_{p}(r)\geq n$, so every $a\equiv r\pmod{q}$ satisfies $p^{n}\mid a$, hence there are no $n$th-power-free integers in the progression for $|a|$ large (so $\mathcal{F}_{n}(X;q,r)=O(1)$).
If $r$ is strictly $n$-admissible, then $v_{p}(r)<\min(\alpha,n)$ for every $p^{\alpha}\parallel q$, so all indicators in the $r$-dependent product vanish, and the main term is independent of $r$. ∎
Dividing (3) by $2X$ and letting $X\to\infty$ shows that the density $\delta_{n}(q,r)$ exists for every residue class $r$, and is given by the constant multiplying $\frac{2X}{q}$. This is exactly what we need later: once a shape (or a union of shapes) is specified by congruence conditions on $a\pmod{M(n)}$, its natural density among $n$th-power-free parameters follows by summing these constants over the relevant residue classes.
Remark 4.9.
Assume $p\parallel n$ and $(a,p)=1$. Proposition 4.1 gives a criterion for the $p$-part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, governed by the congruence $a^{p-1}\equiv 1\pmod{p^{2}}$. Among units modulo $p^{2}$, this congruence holds for a proportion $1/p$. Therefore a positive proportion of parameters $a$ (subject to the $n$th-power-free constraint) satisfy $p\mid[\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, i.e. $\theta$ does not generate $\mathcal{O}_{K_{a}}$. This does not decide whether $K_{a}$ is monogenic via some other generator.
Corollary 4.10.
@@
Definition 4.11 (Shape–admissible residue classes modulo $M(n)$).
A residue class r(modM(n))r\ ($\mathrm${mod}\ M(n)) is called shape–admissible if
$v_{p}(r)\leq 1\qquad\text{for every prime }p\mid n.$
@@
factors through $a\bmod M(n)$. Fix any union $\mathcal{R}\subseteq(\mathbb{Z}/M(n)\mathbb{Z})$ of shape–admissible classes (in the sense of Definition 4.11). Let
@@
Write the indicator of $n$th–power–freeness as $1_{\text{$n$–free}}(a)=\sum_{d^{n}\mid a}\mu(d)$ and count by inclusion–exclusion in the progression $a\equiv r\pmod{M(n)}$. If $r$ is shape–admissible, then $(d,M(n))=1$ in all contributing terms; the main term is
@@
Summing over $r\in\mathcal{R}$ gives the formula; standard partial summation bounds yield $o(X)$.Summing over $r\in\mathcal{R}$ gives the formula; standard partial summation bounds yield $o(X)$.
@@
Here $M(4)=8$ and the shape–admissible residue classes (in the sense after Theorem 4.12) are those r(mod 8)r\ ($\mathrm${mod}\ 8) with $v_{2}(r)\leq 1$, i.e. $r\in\{1,3,5,7,2,6\}$ (six classes). The three quartic shapes are:
@@
By Theorem 4.12, for any union $\mathcal{R}$ of shape–admissible classes one has
@@
(These four densities sum to $\tfrac{24}{36}C_{6}$, reflecting that we have restricted to the 2424 shape–admissible classes.)
@@
$(\mathbb{Z}/p^{t}\mathbb{Z})^{\times}$ is cyclic of order $p^{t-1}(p-1)$. The congruence $x^{p-1}\equiv 1\pmod{p^{t}}$ has exactly p1p-1 solutions. Fix $k\geq 1$. The condition $v_{p}(u^{p-1}-1)\geq k$ is equivalent to $u^{p-1}\equiv 1\pmod{p^{k}}$, so $u\bmod p^{k}$ must lie in one of these p1p-1 solution classes. Each such class modulo $p^{k}$ has $p^{e+1-k}$ lifts to a class modulo $p^{e+1}$, hence the number of units $u\ (\bmod\ p^{e+1})$ with $v_{p}(u^{p-1}-1)\geq k$ is $(p-1)p^{e+1-k}=(p-1)p^{e-(k-1)}$, giving the stated probabilities; the formulae follow. The case $p=2$ is similar with the usual structure of $(\mathbb{Z}/2^{t}\mathbb{Z})^{\times}$. ∎